How much kinetic energy does a paper airplane have if its going 2m/s and has a mass of 0.01 kg? 



1
Expert's answer
2021-01-15T13:46:29-0500

By the the definition of the kinetic energy, we have:


"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot0.01\ kg\cdot(2\ \dfrac{m}{s})^2=0.02\ J."

Answer:

"KE=0.02\ J."